On January 1, 2015, a machine was purchased for $99,900. The machine has an estimated salvage value of $6,660 and an estimated useful life of 5 years. The machine can operate for 111,000 hours before it needs to be replaced. The company closed its books on December 31 and operates the machine as follows: 2015, 22,200 hrs; 2016, 27,750 hrs; 2017, 16,650 hrs; 2018, 33,300 hrs; and 2019, 11,100 hrs.

Part 1

Your answer is correct.

Compute the annual depreciation charges over the machine’s life assuming a December 31 year-end for each of the following depreciation methods. (Round answers to 0 decimal places, e.g. 45,892.)

(1)

Straight-line Method $18,648

(2)

Activity Method

Year

2015 - $18,648

2016 - $23310

2017 – $13, 986

2018 – $27,972

2019 –$ 9,324

(3)

Sum-of-the-Years'-Digits Method

Year

2015 - $31,080

2016 - $24864

2017 – $18,648

2018 – $12, 432

2019 – $6,216

(4)

Double-Declining-Balance Method

Year

2015 - $39,960

2016 - $23,976

2017 - $14,386

2018 –$ 8,631

2019 – $6,287

Part 2

Your answer is incorrect.

Assume a fiscal year-end of September 30. Compute the annual depreciation charges over the asset’s life applying each of the following methods. (Round answers to 0 decimal places, e.g. 45,892.)

Year

Straight-line Method

Sum-of-the-years'-digits method

Double-declining-balance method

2015

$

$

$

2016

2017

2018

2019

2020
Answer #1

Year

Straight line Method

(Note 1)

Sum of Years’-digits method

(Note 2)

Double declining balance

Method(Note 3)

2015

$ 13986

$ 23310

$ 29970

2016

$ 18648

$26418

$ 27972

2017

$ 18648

$20202

$ 16783

2018

$ 18648

$13986

$ 10070

2019

$ 18648

$7770

$ 6042

2020

$ 4662

$ 1554

$ 906

Notes:

  1. Depreciation under Straight Line Method

= (Cost-Residual Value)/number of years

In the first year the machine is used only for 9 months

={ ($99,900- $6,660)/ 5 years}*9 months/12months=18648*9/12=$ 13986

Depreciation for all other 4 years is $ 18648 per year

Depreciation for last year=$ 99900-$6660-$13986-$18648*4=$4662

  1. Depreciation under sum-of-the-years-digits method

Year

Total

Depreciable

cost

Depreciation

Percent

Depreciation

expense

2015

$ 93240

(5/15)*9/12

$ 23310

2016

$ 93240

(3/12*5/15)+(9/12*4/15)

$26418

2017

$ 93240

(3/12*4/15)+(9/12*3/15)

$20202

2018

$ 93240

(3/12*3/15)+(9/12*2/15)

$13986

2019

$ 93240

(3/12*2/15)+(9/12*1/15)

$7770

2020

$ 93240

(3/12*1/15)

$ 1554

  1. Depreciation under Double Declining Method

Rate of depreciation =20%*2=40%

Year

Depreciation Expense

2015

$ 29970

(40%*$99900*9/12)

2016

$ 27972

($99900-$ 29970)0.4

2017

$16783.2 approx $ 16783

($ 69930-$ 27972)*0.4

2018

$ 10069.92 approx$ 10070

($41958-$16783.2)*0.4

2019

$ 6042

($ 25175-$10070)*0.4

2020

$ 906